OpenStudy (anonymous):

Let f(x) = 1/pi(1+x^2), where x ∈ R. a. Show that f is a probability function. b. Find the mean of the random variable X with this probability function. (Note: R is all real numbers.) UPDATE: I've already figured out a), which is 1. However, I still need help on b). Please let me know. Thanks.

OpenStudy (turingtest):

for a probability distribution function\[\int_{-\infty}^{\infty} f(x)dx=1\]

OpenStudy (turingtest):

can you do this indefinite integral?

OpenStudy (anonymous):

I think for a) you can do indefinite, but I don't know about b).

OpenStudy (anonymous):

a. so as turing said if \[\int\limits_{-\infty}^{\infty}f(x)dx=1\] it is a probability density function and the mean will be \[\int\limits_{-\infty}^{\infty}xf(x)dx\]

OpenStudy (turingtest):

I don't why I said indefinite, I meant "improper"

OpenStudy (anonymous):

It should be improper.

OpenStudy (turingtest):

it is when the bounds have an infinity in them lol

OpenStudy (anonymous):

lol, yeah I had to make sure.

OpenStudy (turingtest):

the second integral can be done by parts any problems you are having specifically?

OpenStudy (anonymous):

I think I know what to do with a) now. However, I still don't know what to do for b).

OpenStudy (turingtest):

integration by parts\[dv=\frac1{1+x^2}dx\]\[u=x\]

OpenStudy (turingtest):

though I don't suppose that will converge... :/

OpenStudy (anonymous):

No I don't think it should.